OpenStudy (anonymous):

Subtract 3/x-3 - 4/x+5

OpenStudy (anonymous):

I assume the question is:\[\frac{ 3 }{x-3 }-\frac{ 4 }{x+5 }\]

OpenStudy (anonymous):

\[=\frac{ 3(x+5)-4(x-3) }{ (x-3)(x+5) }\]

OpenStudy (anonymous):

\[=\frac{ 3x+15-4x+12 }{(x-3)(x+5) }\]

OpenStudy (anonymous):

\[=\frac{ 27-x }{(x-3)(x+5) }\]